x+y=3
x-y=13

answer:(8,-5)
need a full break down of how my text book come up with those solutions

Answer by stanbon(75887)

x+y=3
x-y=13
-----------
Add the two equations to get rid of the y-terms:
2x = 16
x = 8
===========
Substitute that into x+y =3
to solve for "Y":
8 + y = 3
y = -5
===============
Cheers,
Stan H.